# Characterizing the nuclear models informed by PREX and CREX: a view from Bayesian inference

### v1

@@ Abstract
New measurements of the weak charge density distributions of 48Ca and 208Pb challenge existing nuclear models. In the post-PREX-CREX era, it is unclear if current models can simultaneously describe weak charge distributions along with accurate measurements of binding energy and charge radii. In this letter, we explore the parameter space of relativistic and non-relativistic models to study the differences between the form factors of the electric and weak charge distributions, $\Delta F=F_{ch}-F_{w}$, in 48Ca and 208Pb. We show, for the first time, the parts of the mean-field models which are the most important in determining the relative magnitude of the neutron skin in lead and calcium nuclei. We carefully disentangle the tension between the PREX/CREX constraints and the ability of the RMF and Skyrme models to accurately describe binding energies and charge radii. We find that the nuclear symmetry energy coefficient $S_{\mathrm{V}}$ and the isovector spin-orbit coefficient $b^{\prime}_{4}$ play different roles in determining $\Delta F$ of 48Ca and 208Pb. Consequently, adjusting $S_{\mathrm{V}}$ or $b^{\prime}_{4}$ shifts predicted $\Delta F$ values toward or away from PREX/CREX measurements. Additionally, $S_{\mathrm{V}}$ and the slope L are marginally correlated given the constraints of our Bayesian inference, allowing us to infer them separately from PREX and CREX data.
@@
Nuclear models describing finite nuclei and nuclear matter – One of the goals in nuclear physics is to construct a unified theory describing the basic properties of a variety of nuclei and neutron star observables. Historically, the binding energy, the charge radii and the response properties of nuclei have been used to calibrate nuclear models. The requirement of simultaneously describing the charge distributions of a variety of isotopes has spurred the improvement of both relativistic mean field (RMF) and non-relativistic Skyrme models, but has remained to be an outstanding challenge for the state-of-the-art nuclear structure theories Lalazissis et al. (1997a); König et al. (2023); Sommer et al. (2022); Malbrunot-Ettenauer et al. (2022); Reinhard and Nazarewicz (2022). It has also been recognized that the neutron distributions in finite nuclei are less sensitive to experimentally measurable quantities like binding energy and charge radii Horowitz and Piekarewicz (2001), leading to less-constrained isovector interactions in traditional nuclear models. Nonetheless, isospin-dependent interactions play a crucial role in describing neutron star properties. Recently, information about differences between the form factors of electric and weak charge distributions $\Delta F=F_{ch}-F_{w}$ of two neutron-rich magic nuclei, 208Pb and 48Ca, has been model-independently measured via parity-violating electron scattering experiments Adhikari et al. (2021a, 2022a).Can existing forms of relativistic RMF models and non-relativistic Skyrme models simultaneously describe the weak charge distributions inferred by PREX and CREX, while maintaining consistency with the basic ground state properties of finite nuclei? If so, what would be the implications for the isospin-dependent interactions of such models? In this letter, we study the parameterizations of existing nuclear models informed by the PREX and the CREX experiments. The linear relationship between the isovector-dependent parts of the nuclear models and the neutron skin thickness of two nuclei (208Pb and 48Ca) discovered in this work for the first time gives the plausible region of $S_{\mathrm{V}}$, $L$ and $b^{\prime}_{4}$ that are mainly constrained by the PREX and CREX, disentangling the information of iso-vector interactions given by neutron skin thickness measurements from those given by the other nuclei property measurements. This linear relationship together with the region of $S_{\mathrm{V}}$, $L$ and $b^{\prime}_{4}$ enhance our understanding of both the nuclei and neutron star properties.
@@
The picture of $S_{\mathrm{v}}-L$ inferred by joint PREX and CREX measurements – For densities of relevance to nuclei, the energy of asymmetric nuclear matter can be expressed as the sum of symmetric nuclear matter (SNM) and symmetry energies as
@@
where $n_{n}$ and $n_{p}$ are the neutron and proton densities, and $n$ is their sum. The SNM energy $E$ can be expanded around the saturation density $n_{0}$:
@@
where $S(n)$ is the density-dependent symmetry energy. An alternative definition based on pure neutron matter (PNM), $S(n)=S(n_{n}=n,n_{p}=0)$, has also been used Gandolfi et al. (2014). The deviation of $S(n)$ between these two definitions represents the breaking of the quadratic approximation, which is usually small for most Skyrme models Lee et al. (1998); Zuo et al. (1999); Drischler et al. (2014). However, such a deviation can be larger in some other models Cai and Chen (2019). In RMF models, the inclusion of the $\delta$-meson breaks the quadratic approximation significantly as shown in Salinas and Piekarewicz (2024). Despite the differences in definitions, $S(n)$ can be expanded around the saturation density as
$S(n)$ $=$ $S_{\mathrm{V}}+\frac{L}{3n_{0}}(n-n_{0})+\frac{K_{sym}}{18n_{0}}(n-n_{0})+...,$ (4)
@@
It has long been recognized that neutron skin thicknesses of neutron-rich nuclei are primarily determined by $S(n)$. The strong correlation between $L$ and the neutron skin thickeness has been extensively studied using Skyrme and RMF models of nuclei Horowitz et al. (2001); Reed et al. (2021). The neutron skin can also be understood using a simple droplet model in terms of the competition between the bulk symmetry term $S_{\mathrm{V}}\cdot I$ and the surface term $\Delta R_{np}\cdot Q$  Myers and Swiatecki (1980); Centelles et al. (2009); Warda et al. (2009). Here, $I=(N-Z)/A$ with $N$ and $Z$ being the neutron and proton numbers in a nucleus and $A$ their sum, $\Delta R_{np}$ is the neutron-proton radius difference, and $Q$ is the surface stiffness parameter. The first of these competing terms is closely related to the properties of bulk matter in the central regions of nuclei and the second relates also to the spin-orbit interaction.
It is not certain that the aforementioned models and approximations work equivalently well to explain the neutron skin thicknesses of all light-to-heavy nuclei. Given the possible sensitivity of neutron skin thickness to both $S_{\mathrm{V}}$ and $L$ (as suggested by previous works) and the weak correlation between them observed in both RMF and Skyrme models, in this work we study the dependence of $\Delta F^{\mathrm{Ca}48}$ and $\Delta F^{\mathrm{Pb}208}$ on $S_{\mathrm{V}}$ and $L$ by extensively surveying the parameter space of both RMF and Skyrme models. Comparing to previous work studying the information content of PREX and CREX Reinhard et al. (2022a); Zhang and Chen (2023); Salinas and Piekarewicz (2023), we apply minimum constraints that are sensitive to isovector interactions except for the PREX and CREX data. This “minimum setting” allows us to characterize the isospin-dependent features of nuclear models mainly informed by PREX and CREX data, rather than the existing extensive measurements of charge radii, binding energy and electric dipole polarizability of finite nuclei.
@@
Figure 1: Symmetry energy parameters $S_{\mathrm{V}}$ and $L$ constrained by parity violating asymmetry of 48Ca (CREX) and 208Pb (PREX). Dashed (dotted) line and the corresponding band represent mean and 1-$\sigma$ band of PREX (CREX) alone based on the correlations in Eq. (9) and Eq. (8). Dash-dotted ellipse represent 1-$\sigma$ contour of joint PREX and CREX constraints. The PREX-2 analysis Reed et al. (2021) as well as constraints from the unitary gas conjecture Tews et al. (2017) and dipole polarizability Piekarewicz et al. (2012) are shown for comparison. The horizontally shaded ellipse show the 1-$\sigma$ contour of the posterior with basic nuclei constraints as the yellow distribution for basic nuclei in Fig. 2, whereas the other two shaded ellipses show the posterior with the +PREX+CREX experiment for Skyrme (vertical shade) and RMF (diagonal shade) models, respectively. The green ellipse shows the UNDEF correlation considering properties of more nuclei Kortelainen et al. (2010). The small black ellipse shows the correlation from chiral EFT calculation of PNMDrischler et al. (2024).
@@
In Fig. 1, we present the constraints of $S_{\mathrm{V}}$ and $L$ given the information of $\Delta F^{\mathrm{Ca}48}$ and/or $\Delta F^{\mathrm{Pb}208}$ without assuming the correlation between them. Given the large number of three-dimensional points of $(\Delta F^{\mathrm{Ca}48},\Delta F^{\mathrm{Pb}208},S_{\mathrm{V}})$ and $(\Delta F^{\mathrm{Ca}48},\Delta F^{\mathrm{Pb}208},L)$ sampled from the posterior distribution of RMF/Skyrme models, we find that $S_{\mathrm{V}}$ and $L$ can be well approximated by a linear combination of $\Delta F^{\mathrm{Ca}48}$ and $\Delta F^{\mathrm{Pb}208}$ as
$S_{\mathrm{V}}$ $=$ $a\Delta F^{Ca48}+b\Delta F^{Pb208}+c$ (5)
$L$ $=$ $a^{\prime}\Delta F^{Ca48}+b^{\prime}\Delta F^{Pb208}+c^{\prime}~{}.$ (6)
@@
The slope coefficients $a$, $b$, $a^{\prime}$, and $b^{\prime}$, as well as the intercept parameters $c$ and $c^{\prime}$, are determined by fitting them to the posterior distribution of $(\Delta F^{Ca48},\Delta F^{Pb208},S_{\mathrm{V}})$ and $(\Delta F^{Ca48},\Delta F^{Pb208},L)$, respectively. These parameters may reflect the inherent structure of RMF and Skyrme models and could be insensitive to the choice of model parameters and the experimental constraints applied in the Bayesian inference. All the sampled points from the posterior that show clear linear relationships are presented in the supplemental material. The fitted slope and intercept parameters are summarized in Table 1. The precision of the fitting formula in Eq. (5) remains within 10% (1-$\sigma$) for both Skyrme and RMF models. Given the mean and standard deviation of $\Delta F^{Pb208}$ and $\Delta F^{Ca48}$ measured by PREX and CREX and the linear relationship in Eq. (5) and Eq. (6), we obtain the mean and covariance of $S_{\mathrm{V}}$ and $L$ to be
@@
in units of MeV for Skyrme and RMF models, respectively. See the red and blue dashdotted ellipses in Fig. 1. The covariance is dominated by the experimental error of $\Delta F^{Ca48}$ or $\Delta F^{Pb208}$. The precision of the fitting formulas in Eq. (5) and Eq. (6) serves as a minor modification.
Based on the linear relationship inferred from the posterior distribution of the chosen RMF and Skyrme models, we further examine the $L$ and $S_{\mathrm{V}}$ values inferred from the PREX and CREX experiments, respectively. Equations (5) and (6) can be rewritten as
$L$ $=$ $\frac{a^{\prime}}{a}S_{\mathrm{V}}-\frac{ba^{\prime}-b^{\prime}a}{a}\Delta F^{\text{Pb208}}-\frac{a^{\prime}c-c^{\prime}a}{a}$ (8)
$L$ $=$ $\frac{b^{\prime}}{b}S_{\mathrm{V}}-\frac{b^{\prime}a-ba^{\prime}}{b}\Delta F^{\text{Ca48}}-\frac{b^{\prime}c-c^{\prime}b}{b}$ (9)
Given the $\Delta F^{\text{Pb208}}$ and $\Delta F^{\text{Ca48}}$ values measured by PREX and CREX, respectively, Eqs. (8) and (9) define a band in the $S_{\mathrm{V}}$-$L$ plane, as shown in Fig. 1. Note that the slope coefficients $a^{\prime}/a$ and $b^{\prime}/b$ are very different. These differing slopes allow us to constrain $S_{\mathrm{V}}$ and $L$ independently using either $\Delta F^{\text{Ca48}}$ or $\Delta F^{\text{Pb208}}$. Our PREX constraint band on $S_{V}-L$ is superior to that of previous PREX-2 analysisReed et al. (2021) as shown in Fig.1, since we disentangle the PREX constraint from the $S_{V}-L$ correlation due to binding energies and charge radii. The overlapping region of these two bands represents the range of ${S_{\mathrm{V}},L}$ inferred by the 1-$\sigma$ experimental measurements from both PREX and CREX, which is quantitatively described by Eq. (7).
The constraints on $S_{\mathrm{V}}$-$L$ correlation outlined above are derived from a combination of the experimental measurements of the neutron skin thickness and the quasi-linear relationships between $\Delta F$ and ${S_{\mathrm{V}},L}$. Notably, these constraints remain independent of the constraints of $S_{\mathrm{V}}$-$L$ based on measurements of nuclei masses and radii, such as those depicted in the UNDEF constraint Kortelainen et al. (2010) shown in Fig. 1. Consequently, one should not anticipate a significant proportion of the sampled points from the posterior distribution of our Bayesian inference to fall within this overlapped region. This is because the Bayesian inference incorporates constraints not only from the PREX and CREX measurements, but also from other fundamental properties of finite nuclei, such as th binding energies and charge radii. Indeed, as illustrated in Fig. 1, the $S_{\mathrm{V}}$ values inferred from measurements of nuclei masses typically range from 30 to 32 MeV, which is in tension with those favored by the combined PREX and CREX measurements.
@@
Figure 2: The posterior distribution of bulk symmetry energy $S_{\mathrm{V}}$ and symmetry energy slope $L$ at saturation density constrained by basic nuclei constraints (dashed yellow), +PREX experiment (dashdotted blue), +CREX experiment (dashdotted red), +PREX+CREX experiment (solid black). Dotted black curve includes additional constraint from dipole Polarizability Piekarewicz et al. (2012). Details of prior and likelihood are discussed in Supplemental Material.
@@
In Fig. 2, we analyze the evolution of the posterior distributions for $S_{\mathrm{V}}$ and $L$ under five different constraint sets: basic nuclear constraints as the ”minimum setting”; inclusion of either PREX or CREX experimental data; inclusion of both PREX and CREX experimental data; and inclusion of additional dipole polarizability constraint. In the upper two panels for the Skyrme model, we observe that the range of $S_{\mathrm{V}}$ is approximately [20,50] MeV, whereas the range of $L$ spans from [0,150] MeV across all Bayesian inferences. The black dot-dashed curve in Fig. 1 illustrates the joint distribution. In the upper right (Skyrme) and lower right (RMF) panels, the posteriors of $L$ exhibit similar trends. The posterior of $L$ constrained by CREX (PREX) measurements has a mean smaller (larger) than that of the posterior constrained by only basic properties of finite nuclei. Moreover, the posterior of $L$ constrained by both PREX and CREX measurements has a mean lower than the one without any constraints from neutron skin measurements, yet slightly higher than the one including only the CREX constraint. This trend is consistent across both our RMF and Skyrme models and aligns with previous studies using Skyrme models Reinhard et al. (2021); Reinhard and Nazarewicz (2022); Zhang and Chen (2023) and RMF models Reed et al. (2021); Salinas and Piekarewicz (2023); Yüksel and Paar (2023). The overlapping integral suggests a probability of inconsistency of 93% (1.8-$\sigma$) for Skyrme models and 76% (1.2-$\sigma$) for RMF models, significantly lower than other analyses showing over 2-$\sigma$ inconsistency Zhang and Chen (2023); Salinas and Piekarewicz (2023); Yüksel and Paar (2023). Further details of our Bayesian inference and Monte Carlo sampling are provided in the supplemental material.
@@
Impact of spin-orbit interactions – Unlike the Hamiltonian of infinite dense matter, the Hamiltonian of finite nuclei incorporates terms proportional to the derivative of isoscalar (vector) densities, primarily arising from spin-orbit interactions. The impact of spin-orbit interactions on single-particle spectra and the charge distributions of various neutron-rich isotopes has been acknowledged and discussed in prior studies Reinhard and Flocard (1995a); Horowitz and Piekarewicz (2012a). However, the isospin-dependent spin-orbit interactions are not well constrained, in comparison with other interactions constrained by isobaric analog states (IAS) Roca-Maza et al. (2018) and mirror nuclei Sagawa et al. (2024). Considering that the values of $F_{\mathrm{ch}}$ and $F_{\mathrm{W}}$ measured in both PREX and CREX are approximately the same, leading to a $\Delta F$ that is roughly one order of magnitude smaller, the influence of spin-orbit terms may play crucial roles in determining the distribution of points of ${\Delta F^{Ca48},\Delta F^{Pb208}}$. For the first time (as far as we are aware), we investigate the influence of spin-orbit interactions on both $\Delta F^{Ca48}$ and $\Delta F^{Pb208}$ across a very large parameter space encompassing both RMF and Skyrme parameterizations. This investigation is based on the posterior distribution incorporating constraints from both PREX and CREX measurements.
@@
Figure 3: Charge and weak from factor difference of 48Ca and 208Pb for samples of RMF (left) and Skyrme (right) model colored by the isovector spin-orbit parameter $b^{\prime}_{4}$ in unit fm4. Colorful lines corresponds to the minimum $\chi^{2}$ fitting of linear function same as Eq.(5) and Eq.(6) with $b^{\prime}_{4}=[-0.2,0,0.2,0.4,0.6,0.8]$. Constraints from unitary gas conjectureTews et al. (2017) and nuclei propertiesKortelainen et al. (2010) calculated from corresponding curves in Fig.1 based on the correlation Eq.(5) and Eq.(6). Blue dashed line shows 1-,2- and 3-$\sigma$ confidence region of joint PREX and CREX results. ‘x’ marker on left panels are RMF models listed in Supplementary material. Reference and $b^{\prime}_{4}$ for Skyrme models with marker ‘x’ is discussed in the text. ‘+’ marker represent the recently calibrated Skyrme modelZhang and Chen (2023), the relativistic Hartree-Bogoliubov modelYüksel and Paar (2023) and RMF model with tensor interactionSalinas and Piekarewicz (2024) including PREX and CREX experiment which take different form of interactions.
@@
$\begin{split}H_{\mathrm{SO}}=b_{4}\textbf{J}\cdot\triangledown n+b^{\prime}_{4}(\textbf{J}_{n}\cdot\triangledown n_{n}+\textbf{J}_{p}\cdot\triangledown n_{p}),\end{split}$ (10)
where the $b_{4}$ and $b^{\prime}_{4}$ coefficients are corresponding to the isoscalar and isovector constributions of the spin-orbit potentials. To compare the influence of spin-orbit interactions in both relativistic and non-relativistic models in a parallel manner, we recast the Dirac equation in Schrödinger form using the Foldy-Wouthuysen transformation and the spin-orbit contributions in RMF models can be approximated as:

### v2

@@ Abstract
New measurements of the weak charge density distributions of 48Ca and 208Pb challenge existing nuclear models. In the post-PREX-CREX era, it is unclear if current models can simultaneously describe weak charge distributions along with accurate measurements of binding energy and charge radii. In this letter, we explore the parameter space of relativistic and non-relativistic models to study the differences between the electric and weak form factors, $\Delta F=F_{ch}-F_{W}$, in 48Ca and 208Pb. We show, for the first time, which aspects of mean-field models are the most important in determining the relative magnitude of the neutron skin in lead and calcium nuclei. We carefully disentangle the tension between the PREX-2/CREX constraints and the ability of the RMF and Skyrme models to accurately describe binding energies and charge radii. We find that the nuclear symmetry energy coefficient $S_{\mathrm{V}}$ and the isovector spin-orbit coefficient $b^{\prime}_{4}$ play different roles in determining $\Delta F$ of 48Ca and 208Pb. Consequently, adjusting $S_{\mathrm{V}}$ or $b^{\prime}_{4}$ shifts predicted $\Delta F$ values toward or away from PREX-2/CREX measurements. Additionally, $S_{\mathrm{V}}$ and the slope L are marginally correlated given the prior constraints of our Bayesian inference, allowing us to infer them separately from PREX-2/CREX data.
@@
Nuclear models describing finite nuclei and nuclear matter – One of the goals in nuclear physics is the construction of a unified theory that describes the basic properties of nuclei and neutron star (NS) observables. Historically, nuclear Hamiltonians have been constrained by binding energies, charge radii, and the energies of giant resonances. The requirement of simultaneously describing the charge distributions of a variety of isotopes has spurred the improvement of nuclear structure theories, including those based on ab initio calculations, Skyrme effective interactions and relativistic mean field (RMF) calculationsLalazissis et al. (1997a); König et al. (2023); Sommer et al. (2022); Malbrunot-Ettenauer et al. (2022); Reinhard and Nazarewicz (2022). The Skyrme force is an effective interaction used in nuclear Hartree-Fock approximations, characterized by a zero range interaction with density-dependent terms. On the other hand, the RMF theory is a relativistic framework describing interacting nucleonic Dirac fields and mesonic mean fields. In RMF theory, the basic fields typically included are the scalar ($\sigma$), vector ($\omega$) and vector-isovector fields. The RMF can be directly related to Skyrme models by considering a non-relativistic expansion Koepf and Ring (1991), which will be discussed in more details in this letter. Neutron distributions in finite nuclei are less sensitive to binding energies and charge radii Horowitz and Piekarewicz (2001), leading to less-constrained isovector interactions in traditional nuclear models. Nonetheless, isospin-dependent interactions play a crucial role in describing NS properties. Recently, information about differences between the electric and weak form factors $\Delta F=F_{ch}-F_{W}$ of two neutron-rich magic nuclei, 208Pb and 48Ca, has been model-independently measured via parity-violating electron scattering experiments Adhikari et al. (2021a, 2022a) and many theoretical models Reed et al. (2024, 2021); Reinhard et al. (2022a); Hagen et al. (2015); Hu et al. (2022); Mahzoon et al. (2017); Atkinson et al. (2020) have been proposed to interpret the measurements from PREX-2 and CREX. However, it is important to note that these models generally succeed in describing either PREX-2 or CREX measurements, but face difficulties in simultaneously explaining both sets of data.
Can existing forms of relativistic RMF models and non-relativistic Skyrme models simultaneously describe the weak charge distributions inferred by PREX-2/CREX, while maintaining consistency with the basic ground state properties of finite nuclei? If so, what would be the implications for the isospin-dependent interactions of such models? In this letter, we investigate this issue by extensively surveying the parameter space of both RMF and Skyrme models to study their dependence on $S_{\mathrm{V}}$, $L$ and isovector spin-orbit interactions. Using Bayesian inference, we analyze the information provided by PREX-2 and CREX measurements, detailed further in the supplemental material and data repository Zhao . Our approach also incorporates the fundamental ground state properties of $^{48}\text{Ca}$, $^{90}\text{Zr}$, and $^{208}\text{Pb}$, in line with previous studies Reinhard et al. (2022a); Zhang and Chen (2023); Salinas and Piekarewicz (2023). Notably, we focus on the physical properties of the Skyrme and RMF models that are consistent with PREX-2/CREX experiments and quantify the extent to which these models can jointly reconcile PREX-2/CREX data with other existing nuclear measurements.
@@
The picture of $S_{\mathrm{v}}-L$ inferred by joint PREX-2/CREX measurements – For densities of relevance to nuclei, the energy of asymmetric nuclear matter can be expressed as the sum of symmetric nuclear matter (SNM) and symmetry energies as
@@
where $n_{n}$ and $n_{p}$ are the neutron and proton densities, and $n$ is their sum. The SNM energy $E$ can be expanded around the saturation density $n_{0}$ , at which the SNM energy reaches its minimum:
@@
where $S(n)$ is the density-dependent symmetry energy. An alternative definition based on pure neutron matter (PNM), $S(n)=S(n_{n}=n,n_{p}=0)$, has also been used Gandolfi et al. (2014). The deviation of $S(n)$ between these two definitions represents the breaking of the quadratic approximation, which is usually small for most Skyrme models Lee et al. (1998); Zuo et al. (1999); Steiner (2006); Drischler et al. (2014). However, such a deviation can be larger in some other models Cai and Chen (2019). In RMF models, the inclusion of the $\delta$-meson with large Yukawa coupling breaks the quadratic approximation significantly as shown in Salinas and Piekarewicz (2024), and Table II in the supplemental material. Ignoring corrections that are quartic in the isospin asymmetry, $S(n)$ can be expanded around the saturation density as
$S(n)$ $=$ $S_{\mathrm{V}}+\frac{L}{3n_{0}}(n-n_{0})+\frac{K_{sym}}{18n_{0}^{2}}(n-n_{0})^{2}+$..., (4)
@@
Neutron skin thicknesses of neutron-rich nuclei are primarily determined by $S(n)$ Steiner et al. (2005a). The strong correlation between $L$ and the neutron skin thickness has been extensively studied using Skyrme and RMF models of nuclei Horowitz et al. (2001); Reed et al. (2021). The neutron skin can also be understood using a simple droplet model in terms of the competition between the bulk symmetry term $S_{\mathrm{V}}\cdot I$ and the surface term $\Delta R_{np}\cdot Q$  Myers and Swiatecki (1980); Centelles et al. (2009); Warda et al. (2009). Here, I=(NZ)/AI=(N-Z)/A with $N$ and $Z$ being the neutron and proton numbers in a nucleus and $A$ their sum, $\Delta R_{np}$ is the neutron-proton radius difference, and $Q$ is the surface stiffness parameter. The first of these competing terms is closely related to the properties of bulk matter in the central regions of nuclei and the second relates also to the spin-orbit interaction.
@@
Figure 1: Symmetry energy parameters $S_{\mathrm{V}}$ and $L$ constrained by parity violating asymmetry of 48Ca (CREX) and 208Pb (PREX-2). The dashed (dotted) line and the corresponding band represent the mean and 1-$\sigma$ band of PREX-2 (CREX) alone based on the correlations in Eq.  (5) and Eq.  (6). Dash-dotted ellipse represents the 1-$\sigma$ contour of joint PREX-2/CREX constraints. The PREX-2 analysis Reed et al. (2021) as well as constraints from the unitary gas conjecture Tews et al. (2017) and dipole polarizability Piekarewicz et al. (2012) are shown for comparison. The horizontally shaded ellipse shows the 1-$\sigma$ contour of the posterior with basic nuclei constraints, corresponding to the yellow distribution shown in Fig. 2, whereas the other two shaded ellipses show the posterior with +PREX+CREX for Skyrme (vertical shade) and RMF (diagonal shade) models. The magenta ellipse shows the UNEDF correlation considering properties of more nuclei Kortelainen et al. (2010a). The small black ellipse shows the correlation from chiral EFT calculation of PNM Drischler et al. (2024).
@@
In Fig. 1, we present the $S_{\mathrm{V}}$ and $L$ constrained by $\Delta F^{\mathrm{Ca}48}$ and $\Delta F^{\mathrm{Pb}208}$, without assuming the correlation between them. Given $\Delta F^{\mathrm{Ca}48}$, $\Delta F^{\mathrm{Pb}208}$, $S_{\mathrm{V}}$ and L sampled from our Bayesian posterior, we find that $S_{\mathrm{V}}$ and $L$ can be well approximated by a linear combination of $\Delta F^{\mathrm{Ca}48}$ and $\Delta F^{\mathrm{Pb}208}$ as
$S_{\mathrm{V}}$ $=$ $a\Delta F^{\mathrm{Ca}48}+b\Delta F^{\mathrm{Pb}208}+c$ (5)
$L$ $=$ $a^{\prime}\Delta F^{\mathrm{Ca}48}+b^{\prime}\Delta F^{\mathrm{Pb}208}+c^{\prime}~.$ (6)
@@
The slope coefficients $a$, $b$, $a^{\prime}$, and $b^{\prime}$, as well as the intercept parameters $c$ and $c^{\prime}$, are determined by fitting them to the posterior distribution of $(\Delta F^{\mathrm{Ca}48},\Delta F^{\mathrm{Pb}208},S_{\mathrm{V}})$ and $(\Delta F^{\mathrm{Ca}48},\Delta F^{\mathrm{Pb}208},L)$, respectively. These parameters may reflect the inherent structure of RMF and Skyrme models and could be insensitive to the Bayesian constraints. All the sampled points from the posterior are presented in the supplemental material. The fitted slope and intercept parameters are summarized in Table 1. The precision of the fitting formula in Eq. (5) remains within 10% (1-$\sigma$) for both Skyrme and RMF models. Given the mean and standard deviation of $\Delta F^{\mathrm{Pb}208}$ and $\Delta F^{\mathrm{Ca}48}$ measured by PREX-2/CREX and the linear relationship in Eq. (5) and Eq. (6), we obtain the mean and covariance of $S_{\mathrm{V}}$ and $L$ to be
@@
in units of MeV for Skyrme and RMF models, respectively. See the red and blue dash-dotted ellipses in Fig. 1. The covariance is dominated by the experimental error of $\Delta F^{\mathrm{Ca}48}$ or $\Delta F^{\mathrm{Pb}208}$. The precision of the fitting formulas in Eq. (5) and Eq. (6) serves as a minor modification.
Comparing to previous PREX-2 analysis Reed et al. (2021), our constraints on $S_{V}$ and L are less affected by the basic finite nuclei ground state properties. The overlapping region of these two bands represents the range of $S_{\mathrm{V}},L$} inferred by the 1-$\sigma$ experimental measurements from PREX-2/CREX, which is quantitatively described by Eq. (7).
@@
Figure 2: The posterior distribution of bulk symmetry energy $S_{\mathrm{V}}$ and symmetry energy slope $L$ at saturation density constrained by basic nuclei constraints (dashed yellow), +PREX experiment (dash-dotted blue), +CREX experiment (dash-dotted red), +PREX+CREX experiment (solid black). The dotted black curve includes additional constraint from dipole Polarizability Piekarewicz et al. (2012). The vertical lines correspond to $1-\sigma$ asymmetric uncertainty. Details of prior and likelihood are discussed in Supplemental Material.
@@
In Fig. 2, we analyze the evolution of the posterior distributions for $S_{\mathrm{V}}$ and $L$ under five different constraint sets: basic nuclear constraints as the ”minimum setting”, corresponding to horizontally shaded ellipse in Fig. 1; inclusion of either PREX-2 or CREX experimental data; inclusion of both PREX-2/CREX experimental data, corresponding to vertically and diagonally shaded ellipse in Fig. 1; and inclusion of additional dipole polarizability constraint. In the upper two panels for the Skyrme model, we observe that the range of $S_{\mathrm{V}}$ is approximately [20,50] MeV, whereas the range of $L$ spans from [0,150] MeV. In the upper right (Skyrme) and lower right (RMF) panels, the posteriors of $L$ exhibit similar trends. The posteriors of $L$ constrained by CREX (PREX-2) measurements has a mean smaller (larger) than that constrained by only basic properties of finite nuclei. Moreover, the posterior of $L$ constrained by both PREX-2 and CREX measurements have a mean lower than that without any neutron skin constraints, but slightly higher than the one that includes only the CREX constraint. This trend exists in both RMF and Skyrme models and aligns with previous studies using SkyrmeReinhard et al. (2021); Reinhard and Nazarewicz (2022); Zhang and Chen (2023) and RMF modelsReed et al. (2021); Salinas and Piekarewicz (2023); Yüksel and Paar (2023). The confidence level integral suggests a probability of inconsistency of 93% (1.8-$\sigma$) for Skyrme models and 76% (1.2-$\sigma$) for RMF models, significantly lower than other analyses showing over 2-$\sigma$ inconsistency Zhang and Chen (2023); Salinas and Piekarewicz (2023); Yüksel and Paar (2023). Further details of our Bayesian inference and Monte Carlo sampling are provided in the supplemental material.
@@
Impact of spin-orbit interactions – Unlike the Hamiltonian of infinite nuclear matter, the Hamiltonian of finite nuclei incorporates terms proportional to the derivative of isoscalar (isovector) densities, primarily arising from spin-orbit interactions. The impact of spin-orbit interactions on single-particle spectra and the charge distributions of various neutron-rich isotopes has been acknowledged and discussed in prior studies Reinhard and Flocard (1995a); Horowitz and Piekarewicz (2012a). However, the isospin-dependent spin-orbit interactions are not well constrained, in comparison with other interactions constrained by isobaric analog states (IAS) Roca-Maza et al. (2018) and mirror nuclei Sagawa et al. (2024). Since the weak charge form factor $F_{W}$ and the charge form factor $F_{ch}$ for $^{208}\mathrm{Pb}$ and $^{48}\mathrm{Ca}$ are very close in value, the resulting difference $\Delta F=F_{W}-F_{ch}$ is relatively small—roughly one order of magnitude smaller than $F_{\mathrm{W}}$ and $F_{\mathrm{ch}}$. As a result, the influence of spin-orbit interactions becomes crucial in accurately determining the distribution of $\{\Delta F_{{}^{48}\mathrm{Ca}},\Delta F_{{}^{208}\mathrm{Pb}}\}$ in our study. For the first time (as far as we are aware), we investigate the influence of spin-orbit interactions on both $\Delta F^{\mathrm{Ca}48}$ and $\Delta F^{\mathrm{Pb}208}$ across a very large parameter space encompassing both RMF and Skyrme parameterizations. This investigation is based on the posterior distribution incorporating constraints from PREX-2/CREX measurements.
@@
Figure 3: Charge and weak from factor difference of 48Ca and 208Pb for samples of RMF (left) and Skyrme (right) model colored by the isovector spin-orbit parameter $b^{\prime}_{4}$ in unit fm4. Colorful lines corresponds to the minimum $\chi^{2}$ fitting of linear function same as Eq.(5) and Eq.(6) with $b^{\prime}_{4}=[-0.2,0,0.2,0.4,0.6,0.8]$. Constraints from unitary gas conjectureTews et al. (2017) and nuclei propertiesKortelainen et al. (2010a) calculated from corresponding curves in Fig.1 based on the correlation Eq.(5) and Eq.(6). Blue dashed line shows 1-,2- and 3-$\sigma$ confidence region of joint PREX-2/CREX results. ‘x’ marker on the left panels are RMF models listed in Supplementary material. Reference and $b^{\prime}_{4}$ for Skyrme models with marker ‘x’ are discussed in the text. ‘+’ marker represent the recently calibrated Skyrme modelZhang and Chen (2023); Yue et al. (2024), the relativistic Hartree-Bogoliubov modelYüksel and Paar (2023) and RMF model with tensor interactionSalinas and Piekarewicz (2024) which take different form of interactions. For readability only part of RMF and Skyrme models are shown. A more comprehensive analysis of the $b^{\prime}_{4}$ distribution of existing models is presented in the supplemental material.
@@
$\begin{split}H_{\mathrm{SO}}=b_{4}\textbf{J}\cdot{\nabla}n+b^{\prime}_{4}(\textbf{J}_{n}\cdot\nabla n_{n}+\textbf{J}_{p}\cdot\nabla n_{p}),\end{split}$ (8)
where the $b_{4}$ and $b^{\prime}_{4}$ coefficients correspond to the isoscalar and isovector contributions of the spin-orbit potentials. To compare spin-orbit interactions in both relativistic and non-relativistic models in a parallel manner, we recast the Dirac equation in Schrödinger form using the Foldy-Wouthuysen transformation and the RMF spin-orbit contributions models can be approximated as: